OpenStudy (anonymous):

A set of data has the following linear regression equation: = 1 + 3x What does the equation predict for a value of x = 13?

OpenStudy (anonymous):

y = 1 + 3x y = 1 + 3*13 y = 1 + 39 y = 40 Answer D